OpenStudy (anonymous):

What is the equation for a line that passes through the points (-2, -3) and (4, 21)? y = -4x + 11 y = 4x - 11 y = 4x + 5 y = -4x - 5

OpenStudy (anonymous):

ill answer it give me a minute

OpenStudy (anonymous):

21--3=24 4--2=6 so slope is 4 Im pretty sure its c

OpenStudy (anonymous):

Yes, thakns. Mind helping out with 2-3 more?

OpenStudy (anonymous):

but wait for melancholymushrooms answer

OpenStudy (anonymous):

no its ok you already answered

OpenStudy (anonymous):

C was right.

OpenStudy (anonymous):

Find Slope. 9x - 4y = -20 -5 5 9/4 -9/4

OpenStudy (anonymous):

9/4

OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, -4) and (7, 10)? 7/17 2/7 7/2 -15/14

OpenStudy (anonymous):

10--4=14 7-3=4

OpenStudy (anonymous):

7/2

OpenStudy (anonymous):

Thank you. I am giving medals. What is the equation for a line that passes through the points (-2,2) and (4,10) y= -2x + 6 y= 2x - 6 y= -2x + 2 y= 2x + 2

OpenStudy (anonymous):

Thats the last one.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

or d wait

OpenStudy (anonymous):

hold on sorry

OpenStudy (anonymous):

ok b